# SpinMeRound: Consistent multiview Identity Generation Using Diffusion Models

### v1

@@ Abstract
Despite recent progress in diffusion models, generating realistic head portraits from novel viewpoints remains a significant challenge. Most current approaches are constrained to limited angular ranges, predominantly focusing on frontal or near-frontal views. Moreover, although the recent emerging large-scale diffusion models have been proven robust in handling 3D scenes, they underperform on facial data, given their complex structure and the uncanny valley pitfalls. In this paper, we propose SpinMeRound, a diffusion-based approach designed to generate consistent and accurate head portraits from novel viewpoints. By leveraging a number of input views alongside an identity embedding, our method effectively synthesizes diverse viewpoints of a subject whilst robustly maintaining its unique identity features. Through experimentation, we showcase our model’s generation capabilities in 360 head synthesis, while beating current state-of-the-art multiview diffusion models. Project page is at spin-me-round.github.io.
@@ 1 Introduction
Recent advances in deep learning have made significant progress in fundamental computer vision tasks, notably in image and video generation [51, 6]. The introduction of diffusion models [26, 16] has transformed these domains by enabling the generation of high-quality visual content, facilitated by the use of large-scale datasets [51, 5]. Despite these developments, the task of generating consistent and accurate head portraits from a single input image still remains a challenging problem. This difficulty is primarily attributed to the limited availability of comprehensive 3D facial datasets, which constrains the training of models, capable of reliably capturing and reconstructing the detailed structure and variations of human faces.
@@
A common practice for modeling human heads from a single image incorporates the use of a 3D Morphable Model (3DMM) [4, 7, 8, 34], to represent the facial shape along with an appearance model [21, 17, 31]. However, given typical training data and the difficulty in modeling complex hairstyles, these methods focus solely on the facial region and avoid or miss the full head and hair. The seminal work of Neural Radiance Fields (NeRF) [40] led to an explosion of works on neural rendering of scenes that could not easily be modeled with textured meshes. Even more, pairing such implicit representations with generative models led to a wide variety of approaches [2, 41, 9] that pushed the boundaries on facial novel view synthesis, achieving high quality and control. More recently, Panohead [1] first showcased high-quality $360^{\circ}$ head portrait synthesis. However, because of its adaptive camera training scheme, the back-head synthesis typically contains many artifacts [32], and its inversion on “in-the-wild” images is challenging and requires complex fine-tuning  [50].
@@
Recently, diffusion models [16] demonstrated superior performance over GAN-based methods in image generation and have achieved great quality in human generation tasks [51]. However, achieving multiview consistency remains a significant challenge. Despite the lack of accurate 3D datasets, recent advances, such as Score Distillation techniques [45, 67], represent an initial step toward leveraging the 2D generation strengths of diffusion models to construct 3D content without any extra training. However, these approaches are computationally intensive, require intricate constraints, and do not consistently yield photorealistic results. Meanwhile, multiview diffusion architectures, employing video diffusion models as a backbone framework [57, 39], still remain very resource-demanding for generating a single novel view, as they rely on comprehensive camera trajectories to generate coherent central objects. Closer to our work, Zero123 [36] introduced a view-conditioned diffusion model that incorporates view features and camera information into the diffusion process. However, the generated images lack strong multiview consistency and are of low quality, which restricts their performance in photorealistic 3D generation. On the other hand, although DiffPortrait3D [24] enables novel view generation through a diffusion process, it requires a scene-specific fine-tuning step while focusing only on near-front views. Other closely related works are Era3D [33] and Morphable Diffusion [10], which generate fixed camera viewpoints, thus limiting their ability to produce full head portraits. The recently proposed Cat3D [20] presents a promising solution by efficiently integrating a number of input views with specified camera poses to achieve consistent novel viewpoint generation. It integrates 3D attention layers for efficiently sharing common information between all views, along with robust camera pose feature maps. However, Cat3D is limited by its lack of focus on human generation and is currently unavailable as an open-source tool.
@@
In this paper, we present SpinMeRound, a multiview diffusion model designed to generate high-fidelity novel views of a given human face. In addition to facial images, our model also generates the corresponding normals, which are typically available for human data and, as we show, improve the model’s performance and consistency on intricate facial features. Moreover, we show that conditioning the model on an identity embedding and one or more input views of a subject during inference, we can not only sample but also reconstruct multiview consistent images from “in-the-wild” facial images. Our method can accurately synthesize photorealistic head portraits from various angles while preserving essential identity characteristics, which can be used to represent or reconstruct 3D scenes. Given the lack of open-source large-scale multiview head datasets and the problematic nature of the permissions of such datasets, our method is solely trained on synthetic data acquired using Panohead [1], making this work accessible to experiment and build with. Overall, in this paper:
@@
  • We introduce SpinMeRound, a multiview diffusion model conditioned on identity embeddings and a number of views that generates novel perspectives of an input subject and their respective normals.
@@
  • We explore its potential by comparing it with current state-of-the-art multiview diffusion-based methods and showcase superior results in full-head portrait generation.
@@ 2.1 Face Modeling
Extensive research has been dedicated to representing 3D faces through a combination of texture maps and 3D meshes, starting with fundamental 3D Morphable Models (3DMM) [4, 7, 8, 34]. These approaches, however, primarily focus on accurately capturing only the frontal region of the head whilst lacking in integrating finer head details such as hair, wrinkles and wearable items. Dealing with this, recent studies have integrated implicit representations for 3D face modeling. Methods such as  [66, 22, 42] integrate Signed Distance Functions (SDFs) whereas other [9, 2, 18, 63] use Neural Radiance Field (NeRF) models [40] for generating photorealistic results. Levering the powerful triplane representation [9], RTRF [56] generates near frontal views in real-time whilst Panohead [1] introduces an adaptive camera training strategy for enabling full head portrait synthesis. An extension of it is 3DPortraitGAN [61] focusing on all round upper body generation. As this work focuses only on full-head generation, Panohead can effectively be used to acquire synthetic portrait datasets.
@@ 2.2 Diffusion Models
Recently, Diffusion models (DMs) [26] have proven their generative abilities by beating the well-established GANs in image synthesis tasks [16]. The availability of large-scale datasets has revolutionized text-to-image generation [51] and video generation [6, 5] tasks. For human face synthesis, a range of methods have emerged to tackle essential tasks, including 3D avatar creation [59, 64], avatar reenactment [30, 14, 15] and texture map generation [43, 19]. We adopt the concatenation strategy outlined in [43, 19] to generate novel viewpoint images and their corresponding shape normals simultaneously. Based on Stable Diffusion [51], Arc2Face [13] and InstantID [58] generate facial images based on an input subject. Especially, Arc2Face exhibits strong generalization abilities in facial image generation, leveraging an up-sampled subset of WebFace42M [69]. Moreover, Arc2Face introduces a robust identity conditioning mechanism and is integrated into our proposed approach.
@@
DreamFusion [45] introduces the use of Score Distillation Sampling (SDS), incorporating a pre-trained text-to-image diffusion model [51] alongside a NeRF model, to synthesize 3D objects from text prompts. In this way, its authors proved that they can efficiently generate 3D objects despite the lack of large-scale datasets by exploiting the generalization abilities of an image generation network. Although subsequent studies focus on better distillation strategies [67, 48, 46], the approaches mentioned above are time-consuming and require complex balancing and additional constraints [20]. Closer to our work is ID-to-3D [3], a score-distilation approach capable of generating 3D faces, however lacking photorealistic results. The authors of Zero123 [36] introduce a multiview diffusion model conditioned on a reference image and a camera pose. Following that, methods such as One-2-3-45 [35], SyncDreamer [37], Consistent1-to-3 [60] and Cascade-Zero123 [62] further focus on multiview consistency by introducing priors during the denoising process. Other studies such as Zero123++ [54], Era3D [33] and Morphable Diffusion [10] generate fixed viewpoints given the input image, without being able to handle arbitrary views. Moreover, in Cat3D [20], a general multiview diffusion model is introduced, using a powerful camera pose conditioning mechanism while using 3D attention layers. Although it has been a robust method for novel view synthesis, it does not focus on facial novel view synthesis, lacks shape normal generation capabilities and is a closed source framework. Closer to our work, DiffPortrait3d [24] showcases novel view capabilities, given an input facial image focusing only on near-front angles. Additionally, it relies on an image-driven approach for camera conditioning and requires a fine-tuning step for each individual subject.
@@
Recently, current video generation models [23, 6, 5] have proven their ability to generate photorealistic models. Methods such as IM-3D [39], V3D [11] and SV3D [57] integrate an off-the-shelf video diffusion model for generating novel viewpoints of a reference object. However, the video generation step makes these methodologies computationally demanding [20]. Additionally, they are often restricted by the requirement for specific camera trajectories, typically revolving around a central subject. In contrast, our approach concentrates on implementing multiview diffusion networks that handle unordered camera poses.
@@ 3 Method
SpinMeRound incorporates a multiview diffusion model, trained on a 3D facial dataset, capable of generating a full head portrait of an input subject, given their identity embedding and a number of views (Fig. 2). In the following sections, we first describe the proposed diffusion model, its training scheme, and the proposed sampling strategy given a single input view.
3.1 Multiview Diffusion Model
@@
SpinMeRound employs a latent multiview diffusion model alongside a powerful identity conditioning mechanism [44], enabling the generation of novel views of an input scene containing a person. Our approach leverages as conditioning inputs $M\in\{1,3\}$ pairs of sparse views of a subject, their respective shape normals and their associated camera poses. Let $\mathbf{I}_{i}$ denote a picked conditioned view and $\bar{\mathbf{I}}_{i}$ a cropped and aligned version of $\mathbf{I}_{i}$. We utilize a pre-trained face recognition network (ArcFace [13]) $\phi$ to extract the identity embedding vector $\mathbf{w}=\phi(\bar{\mathbf{I}}_{i})\in\mathbb{R}^{512}$ capable of incorporating crucial identity features. Then, we inject the identity information into the diffusion model, following Arc2Face [13]. In particular, a text prompt of “a photo of $<$id$>$ person, is fed to a CLIP text encoder [47] followed by the tokenization step. Simultaneously, the identity vector $\mathbf{w}$ is padded to match the embedding dimension, resulting in $\hat{\mathbf{w}}\in\mathbb{R}^{768}$ and replaces the corresponding $<$id$>$ token. The final input token sequence becomes $\mathbf{s}=\{e_{1},e_{2},e_{3},\hat{\mathbf{w}},e_{5}\}$, where $\hat{\mathbf{w}}$ is the padded version of $\mathbf{w}$, and is fed to the encoder $\mathcal{C}$. Finally, we retrieve the corresponding conditioning vector $\mathbf{c}=\mathcal{C}(\mathbf{s})\in\mathbb{R}^{N\times 768}$, where $N$ is the maximum sequence length.
@@
The proposed latent multiview UNet [52] is similar to the one introduced in Cat3D [20]. Our diffusion model is designed to concurrently generate $P=(M+K)=8$ pairs of facial images and their corresponding normals $\mathcal{N}$, given an input conditioning vector $\mathbf{w}$, a number of $M$ views and the $P$ camera poses. Let $I_{i}^{cond}$ represent the input human face images, $I_{j}^{tgt}$ the target face images from novel viewpoints, and $\mathcal{N}_{i}^{cond}$, $\mathcal{N}_{j}^{tgt}$ their respective shape normals, where $i\in\{1,2,...,M\}$, $j\in\{1,2,...,N\}$ with $M$ being the maximum number of the input conditional views and $N$ the number of target viewpoints. Following [43, 20, 19], we employ a pre-trained AutoEncoder of Stable Diffusion 1.5 (SD 1.5) [51], consisting of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. For each input conditioning image, we obtain the corresponding latent feature maps via the encoder $\mathcal{E}$: $\mathbf{z}_{I_{i}^{cond}}=\mathcal{E}(I_{i}^{cond})\in\mathbb{R}^{4\times 64\times 64}$ and $\mathbf{z}_{\mathcal{N}_{i}^{cond}}=\mathcal{E}(\mathcal{N}_{i}^{cond})\in\mathbb{R}^{4\times 64\times 64}$. These maps are then concatenated channel-wise. Similarly, we apply the same procedure to extract the latent vectors for the target viewpoints. The corresponding camera pose information is incorporated for both conditioning and target viewpoints using the mechanism proposed in Cat3D [20]. For each of the $P$ views, the latent feature maps $\mathbf{z}_{i}$ are concatenated channel-wise with the respective ray representation maps [20, 53] $\mathbf{r}_{i}^{cond},\mathbf{r}_{j}^{tgt}\in\mathbb{R}^{149\times 64\times 64}$, which encode the ray origin and direction. A binary mask $\mathbf{m}\in\{\mathbf{0},\mathbf{1}\}^{1\times 64\times 64}$ is then appended to differentiate between conditional and target latent vectors. Finally, we retrieve the conditioning and target latent feature maps $\bar{\mathbf{z}}_{i}^{cond}=\{\mathbf{z}_{I_{i}^{cond}},\mathbf{z}_{\mathcal{N}_{i}^{cond}}\mathbf{r}_{j}^{cond}\mathbf{m}_{j}^{cond}\}$ and $\bar{\mathbf{z}}_{i}^{tgt}=\{\mathbf{z}_{I_{i}^{tgt}},\mathbf{z}_{\mathcal{N}_{i}^{tgt}}\mathbf{r}_{j}^{tgt},\mathbf{m}_{j}^{tgt}\}$.
@@
We initialize our method using a pre-trained LDM model (Arc2Face [44]) trained on large-scale datasets while adding additional attention layers, connecting the multiple latent feature maps. As in  [20], we integrate 3D attention layers [55] by adding them between the original 2D self-attention layers of the LDM, and we fine-tune all layers of the multiview UNet for improved multiview consistency.

### v2

# SpinMeRound: Consistent Multi-View Identity Generation Using Diffusion Models

@@ Abstract
Despite recent progress in diffusion models, generating realistic head portraits from novel viewpoints remains a significant challenge in computer vision. Most current approaches are constrained to limited angular ranges, predominantly focusing on frontal or near-frontal views. Moreover, although the recent emerging large-scale diffusion models have been proven robust in handling 3D scenes, they underperform on facial data, given their complex structure and the uncanny valley pitfalls. In this paper, we propose SpinMeRound, a diffusion-based approach designed to generate consistent and accurate head portraits from novel viewpoints. By leveraging a number of input views alongside an identity embedding, our method effectively synthesizes diverse viewpoints of a subject whilst robustly maintaining its unique identity features. Through experimentation, we showcase our model’s generation capabilities in full head synthesis, while beating current state-of-the-art multi-view diffusion models. Project page: spin-me-round.github.io.
@@ 1 Introduction
Recent advancements in deep learning have brought significant progress to fundamental computer vision tasks, notably in image and video generation [58, 6]. The introduction of Diffusion Models [29, 17] has transformed these domains by enabling the generation of high-quality visual content, facilitated by the use of large-scale datasets [58, 5]. Despite these developments, the task of generating consistent and accurate head portraits from a single input image still remains a challenging problem. This difficulty is primarily attributed to the limited availability of comprehensive 3D facial datasets, which constrains the training of models capable of reliably capturing and reconstructing the detailed structure and variations of human faces.
@@
A common practice for modeling human heads from a single image, incorporates the use of a 3D Morphable Model (3DMM) [4, 7, 8, 38], to represent the facial shape along with an appearance model [22, 18, 35]. However, given typical training data and the difficulty in modeling complex hairstyles, these methods focus solely on the facial region and avoid or miss the full head and hair. The seminal work of Neural Radiance Fields (NeRF) [44] led to an explosion of works on neural rendering of scenes that could not easily be modeled with textured meshes. In addition, pairing such implicit representations with generative models led to a wide variety of approaches [2, 45, 9] that pushed the boundaries on the synthesis of novel facial views, achieving high quality and control. More recently, Panohead [1] first showcased high-quality $360^{\circ}$ head portrait synthesis. However, because of its adaptive camera training scheme, the back-head synthesis typically contains many artifacts [36], and its inversion on “in-the-wild” images is challenging and requires complex fine-tuning  [57].
@@
Recently, diffusion models [17] demonstrated superior performance over GAN-based methods in image generation and have achieved great quality in human generation tasks [58]. However, achieving multi-view consistency remains a significant challenge. Despite the lack of accurate 3D datasets, recent advances, such as Score Distillation techniques [51, 74], represent an initial step toward leveraging the 2D generation strengths of diffusion models to construct 3D content without extra training. Nevertheless, these approaches are computationally intensive, require intricate constraints, and do not consistently yield photorealistic results. Meanwhile, multi-view diffusion architectures, employing video diffusion models as a backbone framework [65, 43], still remain very resource-demanding for generating a single novel view, as they rely on comprehensive camera trajectories to generate coherent central objects. Closer to our work, Zero123 [40] introduced a view-conditioned diffusion model that incorporates view features and camera information into the diffusion process. However, the generated images lack strong multi-view consistency and are of low quality, which restricts their performance in photorealistic 3D generation. Although DiffPortrait3D [26] enables novel view generation through a diffusion process, it still focuses only on near-front views and utilizes a image-conditioned camera mechanism. Other closely related works are Era3D [37] and Morphable Diffusion [10], which generate fixed camera viewpoints, thus limiting their ability to produce full head portraits. The recently proposed Cat3D [21] presents a promising solution by efficiently integrating a number of input views with specified camera poses to achieve a consistent novel viewpoint generation. It integrates 3D attention layers for efficiently sharing common information between all views, along with robust camera pose feature maps. Still, Cat3D is limited by its lack of focus on human generation and is currently unavailable as an open-source tool.
@@
In this paper, we present SpinMeRound, a multi-view diffusion model designed to generate novel high-fidelity views of a given human face. In addition to facial images, our model also generates the corresponding normals, which are typically available for human data and, as we show, improve the model’s performance and consistency on intricate facial features. Moreover, we show that conditioning the model on an identity embedding and one or more input views of a subject during inference, we can not only sample but also reconstruct multi-view consistent images from “in-the-wild” facial images. Our method can accurately synthesize photorealistic head portraits from various angles while preserving essential identity characteristics, which can be used to represent or reconstruct 3D scenes. Given the lack of open-source large-scale multi-view head datasets and the problematic nature of the permissions of such datasets, our method is solely trained on synthetic data acquired using Panohead [1], making this work accessible to experiment and build with. Overall, in this paper:
@@
  • We introduce SpinMeRound, a multi-view diffusion model conditioned on identity embeddings and a number of views that generates novel perspectives of an input subject and their respective normals.
@@
  • We explore its potential by comparing it with current state-of-the-art multi-view diffusion-based methods and showcase superior results in full-head portrait generation.
@@ 2.1 Face Modeling
Extensive research has been dedicated to representing 3D faces through a combination of texture maps and 3D meshes, starting with fundamental 3D Morphable Models (3DMM) [4, 7, 8, 38]. These approaches, however, primarily focus on accurately capturing only the frontal region of the head whilst lacking in integrating finer head details such as hair, wrinkles and wearable items. Dealing with this, recent studies have integrated implicit representations for 3D face modeling. Methods such as  [73, 24, 46] integrate Signed Distance Functions (SDFs) whereas other [9, 2, 19, 70] use Neural Radiance Field (NeRF) models [44] for generating photorealistic results. Levering the powerful triplane representation [9], RTRF [64] generates near-frontal views in real-time whilst Panohead [1] introduces an adaptive camera training strategy for enabling full head portrait synthesis. An extension of it is 3DPortraitGAN [69] focusing on all-around upper body generation. As this work focuses only on full-head generation, Panohead can effectively be used to acquire synthetic portrait datasets.
@@ 2.2 Diffusion Models
Diffusion models (DMs) [29] have proven their generative abilities by beating the well-established GANs in image synthesis tasks [17]. The availability of large-scale datasets has revolutionized text-to-image generation [58] and video generation [6, 5] tasks. For face synthesis, a range of methods have emerged to tackle essential tasks, including 3D avatar creation [67, 71], avatar reenactment [34, 15, 16] and texture map generation [48, 20]. We adopt the concatenation strategy outlined in [48, 20] to generate novel viewpoint images and their corresponding shape normals simultaneously. Based on Stable Diffusion [58], Arc2Face [14] and InstantID [66] generate facial images given an input subject. Notably, Arc2Face exhibits strong generalization abilities in facial image generation, leveraging an up-sampled subset of WebFace42M [76]. Also, it introduces a robust identity conditioning mechanism, which is integrated into our method.
@@
DreamFusion [51] introduces the use of Score Distillation Sampling (SDS), incorporating a pre-trained text-to-image diffusion model [58] alongside a NeRF model, to synthesize 3D objects from text prompts. In this way, its authors proved that they can efficiently generate 3D objects despite the lack of large-scale datasets; by exploiting the generalization abilities of an image generation network. Although subsequent studies focus on better distillation strategies [74, 55, 53], the aforementioned approaches are time-consuming and require complex balancing and additional constraints [21]. Closer to our work, ID-to-3D [3] and Arc2Avatar [23] use score distillation to generate 3D facial geometry, lacking though photorealistic results. Zero123 [40] introduces a multi-view diffusion model conditioned on a reference image and a camera pose. Following it, methods such as One-2-3-45 [39], SyncDreamer [41], Consistent1-to-3 [68] and Cascade-Zero123 [11] further focus on multi-view consistency by introducing priors during the denoising process. Studies such as Zero123++ [61], Era3D [37] and Morphable Diffusion [10] generate fixed viewpoints given the input image, without being able to handle arbitrary views. In Cat3D [21], a general multi-view diffusion model is introduced, using a powerful camera pose conditioning mechanism while using 3D attention layers. Although it has been a robust method for novel view synthesis, it does not focus on facial novel view synthesis, lacks shape normal generation capabilities and is a closed source framework. DiffPortrait3D [26] showcases novel view capabilities, given an input facial image. Unlike our method, which employs image-free camera conditioning, DiffPortrait3D employs an image-driven approach for the desired output viewpoints and can generate only near-front views. More recent works [63, 52] have explored diffusion-based frameworks for generating controllable and animatable avatars. Concurrently to our approach, Pippo [30] and DiffPortrait360 [27] demonstrated $360^{\circ}$ avatar generation, without, though, being able to sample novel identities.
@@
Recently, video generation models [25, 6, 5] have proven their ability to generate photorealistic models. Methods such as IM-3D [43], V3D [12] and SV3D [65] integrate an off-the-shelf video diffusion model for generating novel viewpoints of a reference object. Still, the video generation step makes these methodologies computationally demanding [21]. Additionally, they are often restricted by the requirement for specific camera trajectories, typically revolving around a central subject. In contrast, our approach focuses on implementing multi-view diffusion networks that handle unordered camera poses.
Figure 3: Samples generated by SpinMeRound on “in-the-wild” images: Given only the input images (shown in a small box at the top center), our method produces high-fidelity images from novel angles, along with corresponding shape normals $\mathcal{N}$.
@@ 3 Method
SpinMeRound incorporates a multi-view diffusion model, trained on a 3D facial dataset, capable of generating a full head portrait of an input subject, given their identity embedding and a number of views (Fig. 2). In the following sections, we first describe the proposed diffusion model, its training scheme, and the proposed sampling strategy given a single input view.
3.1 Multi-View Diffusion Model
@@
SpinMeRound employs a latent multi-view diffusion model alongside a powerful identity conditioning mechanism [49], enabling the generation of novel views of an input scene containing a person. Our approach leverages as conditioning inputs $M\in\{1,3\}$ pairs of sparse views of a subject, their respective shape normals and their associated camera poses. Let $\mathbf{I}_{i}$ denote a picked conditioned view and $\bar{\mathbf{I}}_{i}$ a cropped and aligned version of $\mathbf{I}_{i}$. We utilize a pre-trained face recognition network (ArcFace [14]) $\phi$ to extract the identity embedding vector $\mathbf{w}=\phi(\bar{\mathbf{I}}_{i})\in\mathbb{R}^{512}$ capable of incorporating crucial identity features. Then, we inject the identity information into the diffusion model as proposed in Arc2Face [14]. In particular, a text prompt of “a photo of <<id>> person, is fed to a CLIP text encoder [54] followed by the tokenization step. Followed by padding, the identity embedding vector $\mathbf{w}$ replaces the corresponding <<id>> token resulting in $\mathbf{s}=\{e_{1},e_{2},e_{3},\hat{\mathbf{w}},e_{5}\}$, where $\hat{\mathbf{w}}\in\mathbb{R}^{768}$ is the padded version of $\mathbf{w}$. The vector $\mathbf{s}$ is fed to the encoder $\mathcal{C}$, and finally we retrieve the corresponding conditioning vector $\mathbf{c}=\mathcal{C}(\mathbf{s})\in\mathbb{R}^{N\times 768}$, where $N$ is the maximum sequence length.
@@
The proposed latent multi-view UNet [59] is similar to the one introduced in Cat3D [21]. Our diffusion model is designed to concurrently generate $P=(M+K)=8$ pairs of facial images and their corresponding normals $\mathcal{N}$, given an input conditioning vector $\mathbf{w}$, a number of $M$ views and the $P$ camera poses. Let $I_{i}^{cond}$ represent the input human face images, $I_{j}^{tgt}$ the target face images from novel viewpoints, and $\mathcal{N}_{i}^{cond}$, $\mathcal{N}_{j}^{tgt}$ their respective shape normals, where $i\in\{1,2,...,M\}$, $j\in\{1,2,...,N\}$ with $M$ being the maximum number of the input conditional views and $N$ the number of target viewpoints. Following [48, 21, 20], we employ a pre-trained AutoEncoder of Stable Diffusion 1.5 (SD 1.5) [58], consisting of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. For each input conditioning image, we first obtain the corresponding latent feature maps: $\mathbf{z}_{I_{i}^{cond}}=\mathcal{E}(I_{i}^{cond})\in\mathbb{R}^{4\times 64\times 64}$ and $\mathbf{z}_{\mathcal{N}_{i}^{cond}}=\mathcal{E}(\mathcal{N}_{i}^{cond})\in\mathbb{R}^{4\times 64\times 64}$. These are then concatenated channel-wise. Similarly, we apply the same procedure to extract the latent vectors for the target viewpoints. The corresponding camera pose information is incorporated for both conditioning and target viewpoints using the mechanism proposed in Cat3D [21]. For each of the $P$ views, the latent feature maps $\mathbf{z}_{i}$ are concatenated channel-wise with the respective ray representation maps [21, 60] $\mathbf{r}_{i}^{cond},\mathbf{r}_{j}^{tgt}\in\mathbb{R}^{149\times 64\times 64}$, which encode the ray origin and direction. A binary mask $\mathbf{m}\in\{\mathbf{0},\mathbf{1}\}^{1\times 64\times 64}$ is then appended to differentiate between conditional and target latent vectors. Finally, we retrieve the conditioning and target latent feature maps $\bar{\mathbf{z}}_{i}^{cond}=\{\mathbf{z}_{I_{i}^{cond}},\mathbf{z}_{\mathcal{N}_{i}^{cond}}\mathbf{r}_{j}^{cond}\mathbf{m}_{j}^{cond}\}$ and $\bar{\mathbf{z}}_{i}^{tgt}=\{\mathbf{z}_{I_{i}^{tgt}},\mathbf{z}_{\mathcal{N}_{i}^{tgt}}\mathbf{r}_{j}^{tgt},\mathbf{m}_{j}^{tgt}\}$.
@@
As proposed in video diffusion models [5], we initialize our method using a pre-trained LDM model (Arc2Face [49]) trained on large-scale datasets while adding additional attention layers, connecting the multiple latent feature maps. As in  [21], we integrate 3D attention layers [62] by adding them between the original 2D self-attention layers of the LDM, and we fine-tune all layers of the multi-view UNet for improved multi-view consistency.